OpenStudy (anonymous):

Rewrite y = 3x^2 - 24x + 10 in general form.."Completing the Square"

OpenStudy (anonymous):

y=3(x-4)^2-38

OpenStudy (anonymous):

Can you show me how to get it, though?

OpenStudy (anonymous):

3x^2-24x+10=3(x^2-8x)+10=3(x^2-2*x*4+16)+10-48=3(x-4)^2-38